# Exploring nuclear force with pulsar glitch observation

### v1

@@ 2.4 Neutron Pairing Gap
Based on the local density approximation (LDA), we calculate local neutron pairing gap within a WS cell. In the inner crust, neutrons form a ${}^{1}S_{0}$ superfluid. Using the BCS approximation (Khodel et al., 1996; Balberg & Barnea, 1998; Wang & Shen, 2010) and the phenomenological pairing force, the ${}^{1}S_{0}{}$ neutron pairing gap $\Delta(k)$ at zero temperature is obtained by solving the gap equation:
@@
Using the BCS approximation, Eq. (21) removes all contributions beyond the bare nuclear potential in the interaction kernel. In fact, medium polarization, namely the screening of nuclear medium on pairing, leads to an overall suppression to the pairing gap. The reduction of the pairing gap due to medium polarization can be simply represented by the polarization strength $\beta$ (Donati & Pizzochero, 2006), $\Delta_{n}(\beta)=\Delta_{n}/\beta$. Lombardo & Schulze (2000) have found that $\beta$ is between 2 and 3. In this work, we consider only the two cases of the weak polarization $\beta=2$ and strong polarization $\beta=3$.
@@ 2.5 Nuclear Pinning Force
where $r$ is the distance of point $\boldsymbol{x}$ within this vortex to the vortex axis. $\boldsymbol{e}_{\vartheta}$ is the tangent unit vector. Eq. (22) indicates that as the neutrons get closer to the vortex axis, their velocity and kinetic energy increase sharply. In the class limit ($v\ll c$), the kinetic energy per neutron is calculated by $M_{n}^{\star}v^{2}/2$. Thus the kinetic energy per unit volume is written as,
@@
The transition radius, $R_{\infty}$, in the neutron sea is slightly larger than the coherence length $\xi$. The neutron velocity $v$ is always much less than $c$ for $r>R_{\infty}$ or $r>\xi$, the application of the class limit in Eq. (23) is reasonable.
@@
Now we calculated the total energy cost in the IP and NP configurations. At a given density, we set up a cylindrical container with a radius of $R_{\rm{ws}}$ and a height of $2R_{\rm{ws}}$. The IP configuration is obtained after this container is filled with free neutrons and then a vortex is created with its axis coinciding with the axis of the cylinder; if we place the spherical approximated WS cell at the center of this container, while filling the remaining space with free neutrons and then creating a vortex in the same position as the IP configuration, we get the NP configuration. We label $\rho_{n}^{\rm{IP}}(\rho_{\mathrm{B}};\rho,\theta,z)$ and $\rho_{n}^{\rm{NP}}(\rho_{\mathrm{B}};\rho,\theta,z)$ as the neutron density distributions in the IP and NP configurations. Here we need to transfer the 1D density distribution within the WS cell to the density distribution in cylindrical coordinate. Perform the same operations for $\Delta_{n}$, $M_{n}^{\star}$ and $e_{n}$ as for $\rho_{n}$. The pinning energy is obtained by integrating the local difference between the local energy cost $E_{\rm{cost}}^{\rm{IP}}$ and $E_{\rm{cost}}^{\rm{NP}}$,
@@
The pinning force per unit length $f_{\rm{pin}}$ can be estimated (Seveso et al., 2016) by using the pinning energies and properties of WS cells that we have calculated. An extra parameter presented in the calculation is the vortex length $l$, we choose $l=5000R_{\rm{ws}}$ in this work, and a brief discussion of the effects of different $l$ on the calculations can be found in Sec. 3.3.
@@
Figure 7: Pinning energies (upper panels) and pinning forces (lower panels) as a function of $\rho_{\mathrm{B}}$ for DD-ME2 (left panels), PKDD(middle panels), and NL3 (right panels) with $L_{0}=$30, 40, 60, 80 MeV and $\beta=$ 2 and 3. The core-crust transition densities are depicted by the corresponding dashed lines. We only present the results where $E_{\rm{p}}>$ 0, which correspond to the configurations of vortex pinning on the droplet. The results for $E_{\rm{p}}<$ 0 indicate that the pinning force is much weaker than the pinning force just close to the core-crust transition density.
2.6 Superfluidity model for pulsar glitch
For completeness, we briefly introduce the superfluidity model for pulsar glitch in this section. According to the snowplow scenario (Pizzochero, 2011; Seveso et al., 2012), neutron vortices are assumed to be globally rigid and parallel to the rotation axis of the pulsar, with the ends of the vortices terminating at the interface between the outer and inner crust. The balance between the pinning force $F_{\rm{pin}}$,
@@ 2.6 Superfluidity model for pulsar glitch
Due to the dependence of $\Delta\Omega_{\rm{cr}}$ on $r$, the continuously unpinning and repinning of vortices result in a collective outward radial motion of the vortices. Vortices are gradually accumulated in regions of higher $\Delta\Omega_{\rm{cr}}$, ultimately concentrating in the thin sheet near $\Delta\Omega_{\rm{cr}}^{\rm{max}}=\Delta\Omega_{\rm{cr}}(r_{\mathrm{max}})$, where $\Delta\Omega_{\rm{cr}}$ reaches its maximum value at $r=r_{\mathrm{max}}$. When $\Delta\Omega_{\rm{cr}}^{\rm{max}}$ is reached, a large number of vortices depin and transfer the angular momentum to the normal component in the region where $r_{\mathrm{max}}<r<r_{\mathrm{o}}$, causing a vortex avalanche that triggers the occurrence of a glitch. $r_{\mathrm{o}}$ are the inner and outer boundaries of the angular momentum transfer region, respectively.
For an axial symmetry system, the angular velocity $\Omega_{\rm{s}}$ of the superfluid component is proportional to the number $N(r)$ of vortices enclosed in a cylindrical region with $r$. Hence the number of vortices accumulated in peak of $\Delta\Omega_{\rm{cr}}^{\rm{max}}$ is:
@@ 3.1 Dependences of pinning energy and pinning force on nuclear parameters
Fig. 6 shows the pinning force per unit length calculated by DD-ME2 with $L_{0}=$ 30 and 80 MeV throughout the inner crust. We can see that the lower pinning energy of the IP configuration results in a weak pinning region at the bottom of the inner crust, while the NP configuration creates a strong pinning region near the crust-core transition density. The pinning force per unit length in the weak pinning region is about an order of magnitude smaller than that in the strong pinning region. Based on our calculations, $E_{\rm{p}}$ verse $\rho_{\mathrm{B}}$ is qualitatively independent of effective interactions and polarization strengths, while quantitatively dependent on them.
@@
Figure 8: The condensation energy per particle $E_{\rm{cond}}^{\rm{sp}}$ and kinetic energy per particle $E_{\rm{kin}}^{\rm{sp}}$ as a function of $r$ in the cases of bare neutron mass and Dirac neutron mass. The effective interaction is DD-ME2 with $L_{0}=30$ MeV. The matter distribution is calculated at $\rho_{\mathrm{B}}=$0.00631 fm-3. The polarization is taken as $\beta=2$.
@@
Next, we discuss the dependence of the pinning energy and pinning force per unit length on the effective interaction. In the upper panels of Fig. 7, we observe that the pinning energy shows a slight dependence on the interactions in the isoscalar channel, but a significant dependence on the interaction in the isovector channel. When the polarization strength is unchanged, a higher $L_{0}$ corresponds to a lower pinning energy, and a clear maximum in the pinning energy can be observed before the crust-core transition density. As $L_{0}$ decreases, the pinning energy increases, the maximum of the pinning energy shifts to higher densities, and disappears when $L_{0}$ is below 40 MeV. The pinning energy corresponding to a smaller $L_{0}$ may exceed 100 MeV at the bottom of the inner crust, such extreme large pinning energy is not found in previous works and can be explained by the size of the inhomogeneous structure: On the one hand, due to the strong surface tension at small $L_{0}$, the droplet at the bottom of the inner crust can maintain a defined inhomogeneous structure at large $R_{\rm{d}}$ and $R_{\rm{ws}}$; On the other hand, the large size leads to a large-scale volume integral in Eq. (27), so that the small local energy cost differences between IP and NP configurations can accumulate into a large pinning energy. The dependence of the pinning force per unit length on the effective interaction is similar to that of the pinning energy. However, due to the introduction of size effects in the expression described in Seveso et al. (2016), the compact structure at the bottom of the inner crust leads to a reduction in the pinning force per unit length. Therefore, in contrast to the pinning energy, the maximum of the pinning force per unit length can be observed at small $L_{0}$ and disappears at high $L_{0}$. We also find that the dependence of pinning force per unit length on $L_{0}$ is affected by polarization.
@@
The dependence of the pinning energy and pinning force per unit length on polarization strength is direct, i.e., stronger polarization leads to a reduction in both the pinning energy and pinning force per unit length. Since the condensation energy is proportional to $\Delta^{2}$, strong polarization results in a smaller condensation energy. The balance between the local kinetic energy and the local condensation energy under strong polarization is achieved at a larger $r$, and therefore a larger vortex core is obtained. In the case of strong polarization, the vortex core in the NP configuration is able to enclose the entire droplet. As a result, the difference in the kinetic energy contributions between the IP and NP configurations tends to disappear, and then the pinning energy is lowered. Polarization introduces an additional size parameter, i.e., coherence $\xi$, in the estimation of the pinning force per unit length; the dependence of the pinning force on $L_{0}$ under different polarization may be non-monotonic. For example, the pinning force per unit length calculated with PKDD with $L_{0}=40$ MeV under strong polarization is significantly larger than that calculated with PKDD with $L_{0}=30$ MeV, which is in contrast to the case of weak polarization.
@@
In Eqs. (23) and (24), we have replaced the bare mass and Fermi energy used in the non-relativistic calculations with the Dirac effective mass and the exceeded Fermi energy, respectively, thus the nuclear medium effects are considered in our work. In Fig. 8, we show the single-particle kinetic energy $E_{\rm{kin}}^{\rm{sp}}$ and condensation energy $E_{\rm{cond}}^{\rm{sp}}$ distributions within the WS cell in the cases of with and without nuclear medium effects. The effective interaction, polarization, and density are taken as DD-ME2 with $L_{0}=30$ MeV, $\beta=2$, and 0.00631 fm-3, respectively. We have assumed that the density and pairing gap distributions within the WS cell remain unchanged in both cases. We find that, because of the reduction in nucleon mass, the nuclear medium effects lead to a larger kinetic energy contribution, a smaller condensation energy contribution, and a larger vortex core size. Furthermore, we also calculate the pinning energy at $\rho_{\mathrm{B}}=0.00631$ fm-3 (IP configuration) and $\rho_{\mathrm{B}}=0.062$ fm-3 (NP configuration). For the case without nuclear medium effects, the pinning energies are $-1.20$ MeV and 22.97 MeV, respectively; After the nuclear medium effects are considered, the pinning energies change to $-1.05$ MeV and 19.34 MeV. Thus, the smaller pinning energy is expected for both IP and NP configurations when the nuclear medium effects are considered.
@@
Figure 9: (a) Pinning force $F_{\rm{pin}}$, (b) Magnus force $F_{\rm{m}}^{*}$, and (c) Critical angular momentum lag $\Delta\Omega_{\rm{cr}}$ as the function of the distance of the vortex line to the rotational axis $r$. The black and green lines represent the results for $L_{0}=$ 30 and 80 MeV, while solid and dashed lines corresponding to the polarization strengths $\beta=$ 2 and 3, respectively. The effective interaction in isoscalar channel is DD-ME2. We calculate the structure of 1.4 $M_{\odot}$ NS with $R=13.01$ km for $L_{0}=$ 30 MeV and $R=13.60$ km for $L_{0}=$ 80 MeV. Since we mainly focus on the pinning properties of the inner crust, the curves for $r<8$ km are not shown, and they are similar to Fig. 2 and Fig. 3 in Shang & Li (2021). Regarding $\Delta\Omega_{\rm{cr}}$, we observed a extreme large lag near the outer boundary of the inner crust. Due to the magnitude of this lag being nearly unattainable, we truncated $\Delta\Omega_{\rm{cr}}$ to below 0.06 rad/s in the panel (c).
3.2 Constraint from Vela 2000

### v2

@@ 2.4 Neutron Pairing Gap
Based on LDA, we calculate local neutron pairing gap within a WS cell. In the inner crust, neutrons form a ${}^{1}S_{0}$ superfluid. Using the BCS approximation (Khodel et al., 1996; Balberg & Barnea, 1998; Wang & Shen, 2010) and the phenomenological pairing force, the ${}^{1}S_{0}{}$ neutron pairing gap $\Delta(k)$ at zero temperature is obtained by solving the gap equation:
@@
Using the BCS approximation, Eq. (21) removes all contributions beyond the bare nuclear potential in the interaction kernel. In fact, medium polarization, namely the screening of nuclear medium on pairing, leads to an overall suppression to the pairing gap. The reduction of the pairing gap due to medium polarization can be simply represented by the polarization strength $\beta$ (Donati & Pizzochero, 2006), $\Delta_{n}(\beta)=\Delta_{n}/\beta$. Lombardo & Schulze (2000) have found that $\beta$ is between 2.0 and 3.0. In this work, we consider only the two cases of the weak polarization $\beta=2.0$ and strong polarization $\beta=3.0$.
@@ 2.5 Nuclear Pinning Force
where $r$ is the distance of point $\boldsymbol{x}$ within this vortex to the vortex axis. $\boldsymbol{e}_{\vartheta}$ is the tangent unit vector. Eq. (22) indicates that as the neutrons get closer to the vortex axis, their velocity and kinetic energy increase sharply. In the classical limit ($v\ll c$), the kinetic energy per neutron is calculated by $M_{n}^{\star}v^{2}/2$. Thus the kinetic energy per unit volume is written as,
@@
The transition radius, $R_{\infty}$, in the neutron sea is slightly larger than the coherence length $\xi$. The neutron velocity $v$ is always much less than $c$ for $r>R_{\infty}$ or $r>\xi$, the application of the classical limit in Eq. (23) is reasonable.
@@
Now we calculated the total energy cost in the IP and NP configurations. At a given density, we set up a cylindrical container with a radius of $R_{\rm{ws}}$ and a height of $2R_{\rm{ws}}$. The IP configuration is obtained after this container is filled with free neutrons and then a vortex is created with its axis coinciding with the axis of the cylinder; if we place the spherical approximated WS cell at the center of this container, while filling the remaining space with free neutrons and then creating a vortex in the same position as the IP configuration, we get the NP configuration. We label $\rho_{n}^{\rm{IP}}(\rho_{\mathrm{B}};\rho,\theta,z)$ and $\rho_{n}^{\rm{NP}}(\rho_{\mathrm{B}};\rho,\theta,z)$ as the neutron density distributions in the IP and NP configurations. Here we need to transfer the 1D density distribution within the WS cell to the density distribution in cylindrical coordinate. We perform the same operations for $\Delta_{n}$, $M_{n}^{\star}$ and $e_{n}$ as for $\rho_{n}$. The pinning energy is obtained by integrating the local difference between the local energy cost $E_{\rm{cost}}^{\rm{IP}}$ and $E_{\rm{cost}}^{\rm{NP}}$,
@@
The pinning force per unit length $f_{\rm{pin}}$ can be estimated (Seveso et al., 2016) by using the pinning energies and properties of WS cells that we have calculated. This estimation takes into account displacement of the nuclei and averages over all possible orientations of polycrystalline structure of the lattice with respect to the vortex. An extra parameter presented in the calculation is the vortex length $l$, we choose $l=5000R_{\rm{ws}}$ in this work, and a brief discussion of the effects of different $l$ on the calculations can be found in Sec. 3.3.
@@
Figure 7: Pinning energies (upper panels) and pinning forces (lower panels) as a function of $\rho_{\mathrm{B}}$ for DD-ME2 (left panels), PKDD(middle panels), and NL3 (right panels) with $L_{0}=$30, 40, 60, 80 MeV and $\beta=$ 2.0 and 3.0. The core-crust transition densities are depicted by the corresponding dashed lines. We only present the results where $E_{\rm{p}}>$ 0, which correspond to the configurations of vortex pinning on the droplet. The results for $E_{\rm{p}}<$ 0 indicate that the pinning force is much weaker than the pinning force just close to the core-crust transition density.
2.6 The snowplow model for pulsar glitch
For completeness, we briefly introduce the snowplow model for pulsar glitch in this section. According to the snowplow scenario (Pizzochero, 2011; Seveso et al., 2012), neutron vortices are assumed to be globally rigid and parallel to the rotation axis of the pulsar, with the ends of the vortices terminating at the interface between the outer and inner crust. The balance between the pinning force $F_{\rm{pin}}$,
@@ 2.6 The snowplow model for pulsar glitch
Due to the dependence of $\Delta\Omega_{\rm{cr}}$ on $r$, the continuously unpinning and repinning of vortices result in a collective outward radial motion of the vortices. Vortices are gradually accumulated in regions of higher $\Delta\Omega_{\rm{cr}}$, ultimately concentrating in the thin sheet near $\Delta\Omega_{\rm{cr}}^{\rm{max}}=\Delta\Omega_{\rm{cr}}(r_{\mathrm{max}})$, where $\Delta\Omega_{\rm{cr}}$ reaches its maximum value at $r=r_{\mathrm{max}}$. When $\Delta\Omega_{\rm{cr}}^{\rm{max}}$ is reached, a large number of vortices depin and transfer the angular momentum to the normal component in the region where $r_{\mathrm{max}}<r<r_{\mathrm{o}}$, causing a vortex avalanche that triggers the occurrence of a glitch. $r_{\mathrm{o}}$ and $r_{\mathrm{max}}$ are the inner and outer boundaries of the angular momentum transfer region, respectively.
For an axially symmetric system, the angular velocity $\Omega_{\rm{s}}$ of the superfluid component is proportional to the number $N(r)$ of vortices enclosed in a cylindrical region with radial distance $r$. Hence the number of vortices accumulated in peak of $\Delta\Omega_{\rm{cr}}^{\rm{max}}$ is:
@@ 3.1 Dependences of pinning energy and pinning force on nuclear parameters
Fig. 6 shows the pinning force per unit length calculated by DD-ME2 with $L_{0}=$ 30 and 80 MeV throughout the inner crust. We can see that the lower pinning energy of the IP configuration results in a weak pinning region at the top of the inner crust, while the NP configuration creates a strong pinning region near the crust-core transition density. The pinning force per unit length in the weak pinning region is about an order of magnitude smaller than that in the strong pinning region. Based on our calculations, $E_{\rm{p}}$ versus $\rho_{\mathrm{B}}$ exhibit a similar trend for different effective interactions and polarization strengths, with the specific values depending on these factors.
@@
Note that the transition from IP configuration to NP configuration is also observed in previous semi-classical calculation (Pizzochero et al., 1997), although the pinning strengths in our work are weaker than those in Pizzochero et al. (1997) due to the inclusion of polarization. Quantum calculations suggest that NP configuration is more favorable at the top of the inner crust, while the pinning configuration may depend on the effective interaction and polarization at the bottom of the inner crust (Avogadro et al., 2007; Klausner et al., 2023). The difference between two approaches is that quantum calculation takes into account the shell effects and the much larger vortex radius in NP configuration compared to the semiclassical approach, see discussions also in Avogadro et al. (2008).
@@
Figure 8: The condensation energy per particle $E_{\rm{cond}}^{\rm{sp}}$ and kinetic energy per particle $E_{\rm{kin}}^{\rm{sp}}$ as a function of $r$ in the cases of bare neutron mass and Dirac neutron mass. The effective interaction is DD-ME2 with $L_{0}=30$ MeV. The matter distribution is calculated at $\rho_{\mathrm{B}}=$0.00631 fm-3. The polarization is taken as $\beta=2.0$.
@@
Next, we discuss the dependence of the pinning energy and pinning force per unit length on the effective interaction. In the upper panels of Fig. 7, we observe that the pinning energy shows a slight dependence on the interactions in the isoscalar channel, but a significant dependence on the interaction in the isovector channel. When the polarization strength is unchanged, a higher $L_{0}$ corresponds to a lower pinning energy, and a clear maximum in the pinning energy can be observed before the crust-core transition density. As $L_{0}$ decreases, the pinning energy increases, the maximum of the pinning energy shifts to higher densities, and disappears when $L_{0}$ is below 40 MeV. The pinning energy corresponding to a smaller $L_{0}$ may exceed 100 MeV at the bottom of the inner crust, such extreme large pinning energy is not found in previous works and can be explained by the size of the inhomogeneous structure: On the one hand, due to the strong surface tension at small $L_{0}$, the droplet at the bottom of the inner crust can maintain a defined inhomogeneous structure at large $R_{\rm{d}}$ and $R_{\rm{ws}}$; On the other hand, the large size leads to a large-scale volume integral in Eq. (27), so that the small local energy cost differences between IP and NP configurations can accumulate into a large pinning energy. The dependence of the pinning force per unit length on the effective interaction is similar to that of the pinning energy. However, due to the introduction of size effects in the Eq. (21) of Seveso et al. (2016), the compact structure at the bottom of the inner crust leads to a reduction in the pinning force per unit length. Therefore, in contrast to the pinning energy, the maximum of the pinning force per unit length can be observed at small $L_{0}$ and disappears at high $L_{0}$. We also find that the dependence of pinning force per unit length on $L_{0}$ is affected by polarization.
@@
The dependence of the pinning energy and pinning force per unit length on polarization strength is direct, i.e., stronger polarization leads to a reduction in both the pinning energy and pinning force per unit length. Since the condensation energy is proportional to $\Delta^{2}$, strong polarization results in a smaller condensation energy. The balance between the local kinetic energy and the local condensation energy under strong polarization is achieved at a larger $r$, and therefore a larger vortex core is obtained. In the case of strong polarization, the vortex core in the NP configuration is able to enclose the entire droplet. As a result, the difference in the kinetic energy contributions between the IP and NP configurations tends to disappear, and then the pinning energy is lowered. Polarization introduces an additional size parameter, i.e., coherence length $\xi$, in the estimation of the pinning force per unit length; the dependence of the pinning force on $L_{0}$ under different polarization may be non-monotonic. For example, the pinning force per unit length calculated with PKDD with $L_{0}=40$ MeV under strong polarization is significantly larger than that calculated with PKDD with $L_{0}=30$ MeV, which is in contrast to the case of weak polarization.
@@
It is worth mentioning that, treating vortex avalanches as a state-dependent Poisson process, Melatos & Millhouse (2023) performed a Bayesian analysis of the pulsar glitch rates to obtain the maximum $\Delta\Omega_{\rm{cr}}$ and $f_{\rm{pin}}$. Their estimate of $f_{\rm{pin}}$ is independent of our calculations based on nuclear theory. Our calculations indicate a pinning force that is at least three times smaller than the value obtained by Melatos & Millhouse (2023). Additionally, it should be noted that Melatos & Millhouse (2023) did not consider other state-dependent Poisson processes that do not involve vortex avalanches and nuclear pinning, such as starquake, nor did they account for the vortex pinning in the core and general relativistic effects.
@@
In Eqs. (23) and (24), we have replaced the bare mass and Fermi energy used in the non-relativistic calculations with the Dirac effective mass and the exceeded Fermi energy, respectively, thus the nuclear medium effects are considered in our work. In Fig. 8, we show the single-particle kinetic energy $E_{\rm{kin}}^{\rm{sp}}$ and condensation energy $E_{\rm{cond}}^{\rm{sp}}$ distributions within the WS cell in the cases of with and without nuclear medium effects. The effective interaction, polarization, and density are taken as DD-ME2 with $L_{0}=30$ MeV, $\beta=2.0$, and 0.00631 fm-3, respectively. We have assumed that the density and pairing gap distributions within the WS cell remain unchanged in both cases. We find that, because of the reduction in nucleon mass, the nuclear medium effects lead to a larger kinetic energy contribution, a smaller condensation energy contribution, and a larger vortex core size. Furthermore, we also calculate the pinning energy at $\rho_{\mathrm{B}}=0.00631$ fm-3 (IP configuration) and $\rho_{\mathrm{B}}=0.062$ fm-3 (NP configuration). For the case without nuclear medium effects, the pinning energies are $-1.20$ MeV and 22.97 MeV, respectively; After the nuclear medium effects are considered, the pinning energies change to $-1.05$ MeV and 19.34 MeV. Thus, the smaller pinning energy is expected for both IP and NP configurations when the nuclear medium effects are considered.
@@
Figure 9: (a) Pinning force $F_{\rm{pin}}$, (b) Magnus force $F_{\rm{mag}}^{*}$, and (c) Critical angular momentum lag $\Delta\Omega_{\rm{cr}}$ as the function of the distance of the vortex line to the rotational axis $r$. The black and green lines represent the results for $L_{0}=$ 30 and 80 MeV, while solid and dashed lines corresponding to the polarization strengths $\beta=$ 2.0 and 3.0, respectively. The effective interaction in isoscalar channel is DD-ME2. We calculate the structure of 1.4 $M_{\odot}$ NS with $R=13.01$ km for $L_{0}=$ 30 MeV and $R=13.60$ km for $L_{0}=$ 80 MeV. Since we mainly focus on the pinning properties of the inner crust, the curves for $r<8$ km are not shown, and they are similar to Fig. 2 and Fig. 3 in Shang & Li (2021). Regarding $\Delta\Omega_{\rm{cr}}$, we observed a extreme large lag near the outer boundary of the inner crust. Due to the magnitude of this lag being nearly unattainable, we truncated $\Delta\Omega_{\rm{cr}}$ to below 0.06 rad/s in the panel (c).
3.2 Constraints from the 2000 Vela glitch